# On Edge Coloring of Multigraphs

### v4

@@ 3 Extented Tashkinov Tree (ETT)
(1) $\delta\notin\overline{\varphi}(T_{n+1})$. Assume otherwise that $T_{n+1}$ has a different $(\gamma,\delta)$-exit vertex under $\varphi/P$ than $\varphi$. By (A2) for $T_{n+1}$, there is only one $(\gamma,\delta)$-path intersecting $T_{n+1}$ under $\varphi$, so $T_{n+1}$ has exactly one $(\gamma,\delta)$-exit vertex, say $x$, contained in a $(\gamma,\delta)$-exit path $P_{ex}$. Since $P$ does not intersect $T_{n+1}$ and $T_{n+1}$ is $RE$ finished, the coloring $\varphi^{\prime}=\varphi/P$ is $(T_{n+1},D_{n},\varphi)$-stable and therefore is a $T_{n+1}$ mod $\varphi$ coloring by (A2). So $T_{n+1}$ is a closed ETT under $\varphi^{\prime}$ with $n$-rungs. So similarly by (A2) for $T_{n+1}$ under $\varphi^{\prime}$, $T_{n+1}$ also has exactly one $(\gamma,\delta)$-exit vertex under $\varphi^{\prime}$, say $y$, and $y$ is different from $x$. Thus $y$ a root of a $(\gamma,\delta)$-ear $D$ of $T_{n+1}$ under $\varphi$. Because interchanging $\alpha$ and $\beta$ along $P$ changes the $(\gamma,\delta)$-exit vertex of $T_{n+1}$, $\alpha\neq\gamma$ (by the assumption on $\alpha$), and $\alpha\neq\delta$ by (1), we have (2) $P$ shares edges with both $D$ and $P_{ex}$, $\alpha\notin\{\gamma,\delta\}$, and $\beta\in\{\gamma,\delta\}$. We may assume that (3) $y\prec_{L}x$ along $T_{n+1}$. Consequently, $|\overline{\varphi}(T_{n+1}(x))-\overline{\varphi}(x)|\geq 2$ and $|\overline{\varphi}(T_{n+1}(x))|\geq 4$. Since $P$ is disjoint from $T_{n+1}$, the coloring $\varphi^{\prime}$ is $(T_{n+1},D_{n},\varphi)$-stable and therefore is $T_{n+1}$ mod $\varphi$ by (A2). So $T_{n+1}$ is a $RE$ finished closed ETT under $\varphi^{\prime}$ with the same ladder, $F_{i}$, $S_{i}$, and $\Theta_{i}$ for $0\leq i\leq n$. As $\varphi=\varphi^{\prime}/P$, we just consider $T_{n+1}$ under $\varphi^{\prime}$ as a counter-example with $(\gamma,\delta)$-exit vertex $y$ instead, so we have $y\prec_{L}x$. Consequently, since both ends of the uncolored edge $e$ has at least two missing colors, and $T_{n+1}$ is elementary by (A1), we have $|\overline{\varphi}(T_{n+1}(x))-\overline{\varphi}(x)|\geq 2$ and $|\overline{\varphi}(T_{n+1}(x))|\geq 4$.
@@
If $n=0$, then by definition $h=0$ and $T_{0}=\emptyset$, so $x\notin V(T_{h})$. We now assume that $n\geq 1$, $x\in V(T_{h})$, and $\alpha\in\overline{\varphi}(T_{n+1}(x))-\overline{\varphi}(x)$. Note that $h\geq 1$ as $n\geq 1$ and $T_{h}$ is also $RE$ finished, and $y\in V(T_{h})$ by (3). Moreover, by (3) there are two colors $\alpha^{\prime},\gamma^{\prime}\in\overline{\varphi}(T_{h})$. Now this time since $T_{n+1}$ is closed and $RE$ finished, following (A2), switch colors on edges colored by $\gamma$ with $\gamma^{\prime}$ outside of $T_{n+1}$, and switch and edges colored by $\alpha^{\prime}$ with $\alpha$ outside $T_{n+1}$ if $\alpha\notin\overline{\varphi}(T_{h})$, we have a smaller counter-example to (A3) by considering the $RE$ finished ETT $T_{h}$ instead of $T_{n+1}$.
@@
For a path $P$ with vertices $u,v\in V(P)$, denote by $P(u,v)$ the subpath of $P$ from $u$ to $v$.
@@
Let $n$ be a nonnegative integer. Suppose (A1), (A2), (A2.5), (A3) hold for all ETTs with at most $n$ rungs, and (A3.5) holds for all ETTs with at most $(n-1)$-rungs. Let $T_{n+1}$ be a closed ETT with ladder $T_{0}\subset\dots\subset T_{n}\subset T_{n+1}$. Assume $T_{n+1}$ is $RE$ finished and has initial index $h$, $\delta$ is a defective color of $T_{n+1}$ under $\varphi$, $\gamma\in\overline{\varphi}(T_{n+1})$, $x$ is the $(\gamma,\delta)$-exit vertex of $T_{n+1}$ with exit path $P_{ex}$ and $x\notin V(T_{h})$. Suppose $T_{n+1}$ also has an $(\gamma,\delta)$-ear $D$ with one root $y$ such that $y\prec x$ along $T_{n+1}$. Let $\alpha,\beta$ be two colors such that $\alpha\in\overline{\varphi}(T_{n+1})\setminus\{\gamma\}$ and $\beta$ is either $\gamma$ or $\delta$. For any $(\alpha,\beta)$-path $P$, if $P$ does not intersect $T_{n+1}$, then $P$ does not share vertices with both $P_{ex}$ and $D$.
@@
Let everything be as described as in Lemma 4.4, and suppose the $(\alpha,\beta)$-path $P$ shares vertices with both $P_{ex}$ and $D$. We claim that we may assume: (1) $\gamma\in\overline{\varphi}(x)$, $\alpha\in\overline{\varphi}(T_{n+1}(x))$, and there exists a color $\pi\in\overline{\varphi}(T_{n+1}(x))$ such that $\pi\notin(\{\alpha,\gamma\}\cup S_{h})$. Since $y\prec x$ along $T_{n+1}$, both ends of the uncolored edge $e$ are in $T_{n+1}(x)$. Since $\varphi$ has at least $\Delta(G)+1$ many colors, we have $|\overline{\varphi}(T_{n+1}(x))|\geq 4$. Moreover, when $h\geq 1$, we have $|\overline{\varphi}(T_{1})|\geq 5$ by Lemma 3.2. Furthermore, since $x\notin T_{h}$ by the assumption in the statement of the current lemma, there exists colors $\gamma^{\prime}\in\overline{\varphi}(x)$, $\alpha^{\prime}\in\overline{\varphi}(T_{n+1}(x))$, and a color $\pi\in\overline{\varphi}(T_{n+1}(x))$ such that $\pi\notin(\{\alpha^{\prime},\gamma^{\prime}\}\cup S_{h})$. Now since $T_{n+1}$ is closed and $RE$ finished, following (A2), by switching colors on edges colored by $\alpha$ with $\alpha^{\prime}$ and edges colored by $\gamma^{\prime}$ with $\gamma$ outside of $T_{n+1}$ if necessary, we have as claimed.
Our goal is to apply the following technical Lemma 4.5. Now by (1), $P_{ex}$ is the entire $(\gamma,\delta)$-path. Moreover, since $T_{n+1}$ is $RE$ finished, all the $(\gamma_{h},\delta_{h})$-ear intersecting $T_{h}$ when $h\geq 1$ are contained in $G[V(T_{n+1})]$. As a result, if $\delta=\delta_{h}$ with $h\geq 1$, then $P_{ex}$ is disjoint from any $(\gamma_{h},\delta_{h})$-ear intersecting $T_{h}$ Assume the two ends of $P$ are $u$ and $v$. Since $P$ shares vertices with both $P_{ex}$ and $D$, without loss of generality, we may assume the first shared vertex between $P_{ex}$ and $P$ along $P_{ex}$ from $x$ is $u^{\prime}$, the first shared vertex between $D$ and $P$ along $D$ from $y$ is $v^{\prime}$, and $u^{\prime}$ is closer to $u$ than $v^{\prime}$. So now $P_{1}=P_{ex}(xu^{\prime})\cup P(u^{\prime}u)$ and $P_{2}=D(yv^{\prime})\cup P(v^{\prime}v)$ are two edge disjoint paths from $T_{n+1}(x)$ to $\{u,v\}$ satisfying all conditions $(i)$ and $(ii)$ of Lemma 4.5. Thus applying Lemma 4.5, we have both $(II)$ and $(III)$. However, on one side, since $j\leq n+1$ where $j$ is the smallest index with $x\in T_{j}$, by $(II)$ and (A1), the resulting $T$ is elementary under $\mu$. On the other hand, because $u$ and $v$ are different vertices, $T$ is not elementary by $(III)$. This contradiction finishes the proof of Lemma 4.4. ∎
@@
Let $n$ be a nonnegative integer. Suppose (A1), (A2), (A2.5), (A3) hold for all ETTs with at most $n$ rungs, and (A3.5) holds for all ETTs with at most $(n-1)$-rungs. Let $T_{n+1}$ be a closed ETT with ladder $T_{0}\subset\dots\subset T_{n}\subset T_{n+1}$. Assume $T_{n+1}$ has initial index $h$, $\delta$ is a defective color of $T_{n+1}$ under $\varphi$, and $\alpha\in\overline{\varphi}(x)$ where $x$ is the $(\alpha,\delta)$-exit vertex of $T_{n+1}$ with exit path $P_{ex}$. Moreover, if $\delta=\delta_{h}$ when $h\geq 1$, then $P_{ex}$ does not intersect any $(\gamma_{h},\delta_{h})$-ear of $T_{h}$ under $\varphi$. Let $\pi$ be a color with $\pi\in\overline{\varphi}(z)$ and $z\in V(T_{n+1}(x))$ such that $\pi\notin(\{\alpha\}\cup S_{h})$ ($S_{h}=\emptyset$ when $h=0$). Let $P_{1}$ and $P_{2}$ be two edge disjoint paths from $T_{n+1}(x)$ to a vertex set $V$ outside of $T_{n+1}$ satisfying all of the following conditions:
@@
    $T$ is an ETT with the same $(j-1)$-rungs under $\mu$ as $T_{j}$ under $\varphi$.
@@
Let everything be as described as in Lemma 4.5, and suppose $x\notin V(T_{h})$ with $j$ being the smallest index where $x\in T_{j}$. So $j-1\geq h$. Note that $P_{1}$ and $P_{2}$ contain at least two $\delta$ defective boundary edges of $T_{n+1}$ under $\varphi$. Note that $\delta\notin\overline{\varphi}(T_{n+1})$ as $\delta$ is a defective color of $T_{n+1}$ and $T_{n+1}$ is closed under $\varphi$. So by assumptions of the current lemma, $\gamma_{h}\in\overline{\varphi}(T_{h})$, and the elementariness of $T_{n+1}$ by (A1), we have (1) $x\notin V(T_{h})$, colors $\pi,\alpha$, $\delta$, and $\gamma_{h}$ (when $h\geq 1$) are all different, and $\pi\notin S_{h}$. We will recursively construct a series of colorings $\varphi_{i}$ with $i\geq 1$ until we reach the desired $\mu$. We first let $\varphi_{1}^{\prime}=\varphi/P_{x}(\alpha,\delta,\varphi)$, and then construct $\varphi_{1}$ from $\varphi_{1}^{\prime}$ by switching all the $\alpha$ colored edges of $P_{1}\cup P_{2}$ not in $P_{z}(\pi,\alpha,\varphi_{1}^{\prime})$ to $\pi$ by Kempe changes. Let $T^{1}$ be a closure of $T_{j}(x)$ under $\varphi_{1}$. We claim that $\varphi_{1}$ satisfies the following:
@@
    $T^{1}$ is an elementary ETT with the same $j-1$-rungs under $\varphi_{1}$ as $T_{j}$ under $\varphi$.
@@
    All the $\alpha$-colored edges of $P_{1}\cup P_{2}$ under $\varphi_{1}$ belong to $P_{z}(\pi,\alpha,\varphi_{1})$ after the vertex $x$ along the direction from $z$.
@@
Clearly $\varphi_{1}^{\prime}=\varphi/P_{x}(\alpha,\delta,\varphi)$ is $(T_{j}(x)-x,D_{j-1},\varphi)$-stable, because $P_{x}(\alpha,\delta,\varphi)\cap T_{n+1}=\{x\}$. As $T^{i-1}$ is closed for both $\alpha_{i}$ and $\pi$ under $\varphi$, all the $(\pi,\alpha)$-chains intersect $T_{n+1}$ under $\varphi_{1}^{\prime}$ are contained in $G[V(T_{n+1})]$ except for $P_{z}(\pi,\alpha,\varphi_{1}^{\prime})$ which contains $x$ who is the only $(\pi,\alpha)$-exit of $T_{n+1}$ under $\varphi_{1}^{\prime}$, and all vertices in $P_{z}(\pi,\alpha,\varphi_{1}^{\prime})$ from $z$ to $x$ are contained in $T_{n+1}$. Moreover, by (i) in the statement of the current lemma, (1), and $P_{x}(\alpha,\delta,\varphi)\cap T_{n+1}=\{x\}$, we see that $\pi$ is not used on edges of $P_{1}\cup P_{2}-G[V(T_{n+1})]$ and $\alpha$ is not used on edges of $(P_{1}\cup P_{2})\cap G[V(T_{n+1})]$ under $\varphi_{1}^{\prime}$. Therefore, since all vertices in $P_{z}(\pi,\alpha,\varphi_{1}^{\prime})$ from $z$ to $x$ are contained in $T_{n+1}$, none of the $\alpha$ colored edges of $P_{1}\cup P_{2}$ under $\varphi_{1}^{\prime}$ is in the subpath of $P_{z}(\pi,\alpha,\varphi_{1}^{\prime})$ from $z$ to $x$. Consequently, we have $(e_{1})$ because we switch all the $\alpha$ colored edges of $P_{1}\cup P_{2}$ not in $P_{z}(\pi,\alpha,\varphi_{1}^{\prime})$ to $\pi$ by Kempe changes, and none of these $(\pi,\alpha)$-chains intersects $T_{n+1}$ under $\varphi_{1}^{\prime}$. It also follows that
@@
By $(a_{1})$ and (2), we see that $T_{h}$ has the same $(\gamma_{h},\delta_{h})$-exit vertex under $\varphi_{1}$ as under $\varphi$ if $h\geq 1$. Since $\varphi_{1}$ is $(T_{j}(x)-x,D_{j-1},\varphi)$-stable and $x\notin T_{j-1}$, we see that $\varphi_{1}$ is also $(T_{j-1},D_{j-1},\varphi)$-stable, and thus is $T_{j-1}+f_{j-1}$ mod $\varphi$ following (A2.5), and therefore $T_{j}(x)$ is an ETT with $(j-1)$-rungs under $\varphi_{1}$ by Corollary 3.1. Since $T^{1}$ is a closure of $T_{j}(x)$ under $\varphi_{1}$ and $j-1\leq n$, we have $(b_{1})$ by (A1) on $T^{1}$.

### v5

@@ 3 Extented Tashkinov Tree (ETT)
(1) $\delta\notin\overline{\varphi}(T_{n+1})$.
Assume otherwise that $T_{n+1}$ has a different $(\gamma,\delta)$-exit vertex under $\varphi/P$ than $\varphi$. By (A2) for $T_{n+1}$, there is only one $(\gamma,\delta)$-path intersecting $T_{n+1}$ under $\varphi$, so $T_{n+1}$ has exactly one $(\gamma,\delta)$-exit vertex, say $x$, contained in a $(\gamma,\delta)$-exit path $P_{ex}$. Since $P$ does not intersect $T_{n+1}$ and $T_{n+1}$ is $\operatorname{RE}$ finished, the coloring $\varphi^{\prime}=\varphi/P$ is $(T_{n+1},D_{n},\varphi)$-stable and therefore is a $T_{n+1}$ mod $\varphi$ coloring by (A2). So $T_{n+1}$ is a closed ETT under $\varphi^{\prime}$ with $n$-rungs. Therefore, by (A2) for $T_{n+1}$ under $\varphi^{\prime}$, $T_{n+1}$ also has exactly one $(\gamma,\delta)$-exit vertex under $\varphi^{\prime}$, say $y$, and $y$ is different from $x$. Thus $y$ a root of a $(\gamma,\delta)$-ear $D$ of $T_{n+1}$ under $\varphi$. Because interchanging $\alpha$ and $\beta$ along $P$ changes the $(\gamma,\delta)$-exit vertex of $T_{n+1}$, $\alpha\neq\gamma$ (by the assumption on $\alpha$), and $\alpha\neq\delta$ by (1), we have
(2) $P$ shares edges with both $D$ and $P_{ex}$, $\alpha\notin\{\gamma,\delta\}$, and $\beta\in\{\gamma,\delta\}$.
We may assume that
(3) $y\prec x$ along $T_{n+1}$. Consequently, $|\overline{\varphi}(T_{n+1}(x))-\overline{\varphi}(x)|\geq 2$ and $|\overline{\varphi}(T_{n+1}(x))|\geq 4$.
Since $P$ is disjoint from $T_{n+1}$, the coloring $\varphi^{\prime}$ is $(T_{n+1},D_{n},\varphi)$-stable and therefore is $T_{n+1}$ mod $\varphi$ by (A2). So $T_{n+1}$ is a $\operatorname{RE}$ finished closed ETT under $\varphi^{\prime}$ with the same ladder, $F_{i}$, $S_{i}$, and $\Theta_{i}$ for $0\leq i\leq n$. As $\varphi=\varphi^{\prime}/P$, we just consider $T_{n+1}$ under $\varphi^{\prime}$ as a counterexample with $(\gamma,\delta)$-exit vertex $y$ instead, so we have $y\prec x$. Consequently, since both ends of the uncolored edge $e$ have at least two missing colors, and $T_{n+1}$ is elementary by (A1), we have $|\overline{\varphi}(T_{n+1}(x))-\overline{\varphi}(x)|\geq 2$ and $|\overline{\varphi}(T_{n+1}(x))|\geq 4$.
@@
If $n=0$, then by definition $h=0$ and $T_{0}=\emptyset$, so $x\notin V(T_{h})$. We then consider the case that $n\geq 1$, $x\in V(T_{h})$, and $\alpha\in\overline{\varphi}(T_{n+1}(x))-\overline{\varphi}(x)$. Note that $h\geq 1$ as $n\geq 1$ and $T_{h}$ is also $\operatorname{RE}$ finished, and $y\in V(T_{h})$ by (3). Moreover, by (3) there are two colors $\alpha^{\prime},\gamma^{\prime}\in\overline{\varphi}(T_{h})$. Now since $T_{n+1}$ is closed and $\operatorname{RE}$ finished, following (A2), by first switching colors on edges colored by $\gamma$ with $\gamma^{\prime}$ outside of $T_{n+1}$, and then switching the edges colored by $\alpha^{\prime}$ with $\alpha$ outside $T_{n+1}$ if $\alpha\notin\overline{\varphi}(T_{h})$, we have a smaller counterexample to (A3) by considering the $\operatorname{RE}$ finished ETT $T_{h}$ instead of $T_{n+1}$.
@@
For a path $P$ with vertices $u,v\in V(P)$, denote by $P[u,v]$ the subpath of $P$ from $u$ to $v$.
@@
Let $n$ be a nonnegative integer. Suppose (A1), (A2), (A2.5), (A3) hold for all ETTs with at most $n$ rungs, and (A3.5) holds for all ETTs with at most (n1)(n-1)-rungs. Let $T_{n+1}$ be a closed ETT with ladder $T_{0}\subset\dots\subset T_{n}\subset T_{n+1}$. Assume $T_{n+1}$ is $\operatorname{RE}$ finished and has initial index $h$, $\delta$ is a defective color of $T_{n+1}$ under $\varphi$, $\gamma\in\overline{\varphi}(T_{n+1})$, $x$ is the $(\gamma,\delta)$-exit vertex of $T_{n+1}$ with exit path $P_{ex}$ and $x\notin V(T_{h})$. Suppose $T_{n+1}$ also has an $(\gamma,\delta)$-ear $D$ with one root $y$ such that $y\prec x$ along $T_{n+1}$. Let $\alpha,\beta$ be two colors such that $\alpha\in\overline{\varphi}(T_{n+1})\setminus\{\gamma\}$ and $\beta$ is either $\gamma$ or $\delta$. For any $(\alpha,\beta)$-path $P$, if $P$ does not intersect $T_{n+1}$, then $P$ does not share vertices with both $P_{ex}$ and $D$.
@@
Let everything be as described as in Lemma 4.4, and suppose the $(\alpha,\beta)$-path $P$ shares vertices with both $P_{ex}$ and $D$. We claim that we may assume:
(1) $\gamma\in\overline{\varphi}(x)$, $\alpha\in\overline{\varphi}(T_{n+1}(x))$, and there exists a color $\pi\in\overline{\varphi}(T_{n+1}(x))$ such that $\pi\notin(\{\alpha,\gamma\}\cup S_{h})$.
Since $y\prec x$ along $T_{n+1}$, both ends of the uncolored edge $e$ are in $T_{n+1}(x)$. Since $\varphi$ has at least $\Delta(G)+1$ many colors, we have $|\overline{\varphi}(T_{n+1}(x))|\geq 4$. Moreover, when $h\geq 1$, we have $|\overline{\varphi}(T_{1})|\geq 5$ by Lemma 3.2. Furthermore, since $x\notin T_{h}$ by the assumption in the statement of the current lemma, there exist colors $\gamma^{\prime}\in\overline{\varphi}(x)$, $\alpha^{\prime}\in\overline{\varphi}(T_{n+1}(x))$, and a color $\pi\in\overline{\varphi}(T_{n+1}(x))$ such that $\pi\notin(\{\alpha^{\prime},\gamma^{\prime}\}\cup S_{h})$. Now since $T_{n+1}$ is closed and $\operatorname{RE}$ finished, following (A2), by switching colors on edges colored by $\alpha$ with $\alpha^{\prime}$ and edges colored by $\gamma^{\prime}$ with $\gamma$ outside of $T_{n+1}$ if necessary, we have as claimed.
Our goal is to apply the following technical Lemma 4.5 with $\gamma$ in place of $\alpha$. Now by (1), $P_{ex}$ is the entire $(\gamma,\delta)$-path. Moreover, since $T_{n+1}$ is $\operatorname{RE}$ finished, all the $(\gamma_{h},\delta_{h})$-ear intersecting $T_{h}$ when $h\geq 1$ are contained in $G[V(T_{n+1})]$. As a result, if $\delta=\delta_{h}$ with $h\geq 1$, then $P_{ex}$ is disjoint from any $(\gamma_{h},\delta_{h})$-ear intersecting $T_{h}$ . Assume the two ends of $P$ are $u$ and $v$. Since $P$ shares vertices with both $P_{ex}$ and $D$, without loss of generality, we may assume the first shared vertex between $P_{ex}$ and $P$ along $P_{ex}$ from $x$ is $u^{\prime}$, the first shared vertex between $D$ and $P$ along $D$ from $y$ is $v^{\prime}$, and $u^{\prime}$ is closer to $u$ than $v^{\prime}$ along $P$. So now $P_{1}=P_{ex}[x,u^{\prime}]\cup P[u^{\prime},u]$ and $P_{2}=D[y,v^{\prime}]\cup P[v^{\prime},v]$ are two edge disjoint paths from $T_{n+1}(x)$ to $\{u,v\}$ satisfying all conditions $(i)$ and $(ii)$ of Lemma 4.5. Thus applying Lemma 4.5, we have both $(II)$ and $(III)$. However, on one side, since $j\leq n+1$ where $j$ is the smallest index with $x\in T_{j}$, by $(II)$ and (A1), the resulting $T$ is elementary under $\mu$. On the other hand, because $u$ and $v$ are different vertices, $T$ is not elementary by $(III)$. This contradiction finishes the proof of Lemma 4.4. ∎
@@
Let $n$ be a nonnegative integer. Suppose (A1), (A2), (A2.5), (A3) hold for all ETTs with at most $n$ rungs, and (A3.5) holds for all ETTs with at most (n1)(n-1)-rungs. Let $T_{n+1}$ be a closed ETT with ladder $T_{0}\subset\dots\subset T_{n}\subset T_{n+1}$. Assume $T_{n+1}$ has initial index $h$, $\delta$ is a defective color of $T_{n+1}$ under $\varphi$, and $\alpha\in\overline{\varphi}(x)$ where $x$ is the $(\alpha,\delta)$-exit vertex of $T_{n+1}$ with exit path $P_{ex}$. Moreover, if $\delta=\delta_{h}$ when $h\geq 1$, then $P_{ex}$ does not intersect any $(\gamma_{h},\delta_{h})$-ear of $T_{h}$ under $\varphi$. Let $\pi$ be a color with $\pi\in\overline{\varphi}(z)$ and $z\in V(T_{n+1}(x))$ such that $\pi\notin(\{\alpha\}\cup S_{h})$ ($S_{h}=\emptyset$ when $h=0$). Let $P_{1}$ and $P_{2}$ be two edge disjoint paths from $T_{n+1}(x)$ to a vertex set $V$ outside of $T_{n+1}$ satisfying all of the following conditions:
@@
    $T$ is an ETT with the same (j1)(j-1)-rungs under $\mu$ as $T_{j}$ under $\varphi$.
@@
Let everything be as described as in Lemma 4.5, and suppose $x\notin V(T_{h})$ with $j$ being the smallest index where $x\in T_{j}$. So $j-1\geq h$. Note that $P_{1}$ and $P_{2}$ together contain at least two $\delta$ defective boundary edges of $T_{n+1}$ under $\varphi$. Moreover, $\delta\notin\overline{\varphi}(T_{n+1})$ as $\delta$ is a defective color of $T_{n+1}$ and $T_{n+1}$ is closed under $\varphi$. So by assumptions of the current lemma, $\gamma_{h}\in\overline{\varphi}(T_{h})$, and the elementariness of $T_{n+1}$ by (A1), we have
(1) $x\notin V(T_{h})$, colors $\pi,\alpha$, $\delta$, and $\gamma_{h}$ (when $h\geq 1$) are all different, and $\pi\notin S_{h}$.
We will recursively construct a series of colorings $\varphi_{i}$ with $i\geq 1$ until we reach the desired $\mu$. We first let $\varphi_{1}^{\prime}=\varphi/P_{x}(\alpha,\delta,\varphi)$, and then construct $\varphi_{1}$ from $\varphi_{1}^{\prime}$ by switching all the $\alpha$ colored edges of $P_{1}\cup P_{2}$ not in $P_{z}(\pi,\alpha,\varphi_{1}^{\prime})$ to $\pi$ by Kempe changes. Let $T^{1}$ be a closure of $T_{j}(x)$ under $\varphi_{1}$. We claim that $\varphi_{1}$ satisfies the following:
@@
    $T^{1}$ is an elementary ETT with the same (j1)(j-1)-rungs under $\varphi_{1}$ as $T_{j}$ under $\varphi$.
@@
    All the $\alpha$ colored edges of $P_{1}\cup P_{2}$ under $\varphi_{1}$ belong to $P_{z}(\pi,\alpha,\varphi_{1})$ after the vertex $x$ along the direction from $z$.
@@
Clearly $\varphi_{1}^{\prime}=\varphi/P_{x}(\alpha,\delta,\varphi)$ is $(T_{j}(x)-x,D_{j-1},\varphi)$-stable, because $P_{x}(\alpha,\delta,\varphi)\cap T_{n+1}=\{x\}$. As $T_{n+1}$ is closed for both $\alpha$ and $\pi$ under $\varphi$, all the $(\pi,\alpha)$-chains intersecting $T_{n+1}$ under $\varphi_{1}^{\prime}$ are contained in $G[V(T_{n+1})]$ except for $P_{z}(\pi,\alpha,\varphi_{1}^{\prime})$ which contains $x$ who is the only $(\pi,\alpha)$-exit of $T_{n+1}$ under $\varphi_{1}^{\prime}$, and all vertices in $P_{z}(\pi,\alpha,\varphi_{1}^{\prime})$ from $z$ to $x$ are contained in $T_{n+1}$. Moreover, by (i) in the statement of the current lemma, (1), and $P_{x}(\alpha,\delta,\varphi)\cap T_{n+1}=\{x\}$, we see that $\pi$ is not used on edges of $P_{1}\cup P_{2}-G[V(T_{n+1})]$ and $\alpha$ is not used on edges of $(P_{1}\cup P_{2})\cap G[V(T_{n+1})]$ under $\varphi_{1}^{\prime}$. Therefore, since all vertices in $P_{z}(\pi,\alpha,\varphi_{1}^{\prime})$ from $z$ to $x$ are contained in $T_{n+1}$, none of the $\alpha$ colored edges of $P_{1}\cup P_{2}$ under $\varphi_{1}^{\prime}$ is in the subpath of $P_{z}(\pi,\alpha,\varphi_{1}^{\prime})$ from $z$ to $x$. Consequently, we have $(e_{1})$ because we switch all the $\alpha$ colored edges of $P_{1}\cup P_{2}$ not in $P_{z}(\pi,\alpha,\varphi_{1}^{\prime})$ to $\pi$ by Kempe changes, and none of these $(\pi,\alpha)$-chains intersects $T_{n+1}$ under $\varphi_{1}^{\prime}$. It also follows that
@@
By $(a_{1})$ and (2), we see that $T_{h}$ has the same $(\gamma_{h},\delta_{h})$-exit vertex under $\varphi_{1}$ as under $\varphi$ if $h\geq 1$. Since $\varphi_{1}$ is $(T_{j}(x)-x,D_{j-1},\varphi)$-stable and $x\notin T_{j-1}$, we see that $\varphi_{1}$ is also $(T_{j-1},D_{j-1},\varphi)$-stable, and thus is $T_{j-1}+f_{j-1}$ mod $\varphi$ following (A2.5), and therefore $T_{j}(x)$ is an ETT with (j1)(j-1)-rungs under $\varphi_{1}$ by Corollary 3.1. Since $T^{1}$ is a closure of $T_{j}(x)$ under $\varphi_{1}$ and $j-1\leq n$, we have $(b_{1})$ by (A1) on $T^{1}$.